OpenStudy (anonymous):

Please help with this.. NONLINEAR ODEs Using a method of this section or separating variables,findthe general solution. If an initial condition is given,find also the particular solution and sketch or graph it. y'=3.2y-10y^2

OpenStudy (loser66):

What are we supposed to do?

OpenStudy (loser66):

first off, you have to have "given condition" then, just replace y' = dy/dx, \(\dfrac{dy}{3.2y -10y^2}=dx\) then take integral both sides, then plug the given condition on to find C Plug that C back to get the whole answer. I took the course long time ago, not sure whether this logic works or not. just try. :)